OpenStudy (judson_rockets):

The surface area of a cube is 600 cm2. What is the area of one face of this cube? cm2

OpenStudy (abdullahm):

Do you know the formula for surface area of a cube?

OpenStudy (judson_rockets):

NOPE

OpenStudy (abdullahm):

Surface area of a cube = \(\bf 6a^2\)

OpenStudy (abdullahm):

where a is the side length. And the question wants us to find the area of one face of the cube which is a * a = a^2

OpenStudy (abdullahm):

They gave us the surface area as 600 So solve for a^2 \(\bf 600 = 6a^2\) Just divide both sides by 6 to get your final answer. :)

OpenStudy (judson_rockets):

So 100

OpenStudy (abdullahm):

Correct!

OpenStudy (aziaboo45):

yeah